Last year, you purchased 660 shares of Forever, Inc., stock at a price of $48.74 per share. You received $990 in dividends and a total of $38,353 when you sold the stock. What was the capital gains yield on this stock?
Answer #1

Sale Price = 38,353 / 660 = 58.11

Capital gain Yield = P1 / P0 - 1 = 58.11 / 48.74 - 1 = 19.23%